an acceleration force of 500N changes an object from 40m/sec to 60m/sec in a distance of 100m. the mass is 25 kg. determine the force of friction acting opposite this force
1
Expert's answer
2017-10-25T15:10:07-0400
The acceleration is
a=(v^2-v_0^2)/2s=(60^2 - 40^2)/(2⋅100)=10(m⋅s^(-2) ).

From Second Newton's law we obtain:F-F_fr=ma.

Thus, the friction force is F_fr=F-ma=500-25⋅10=250(N).